A club would like to know what the learning rate is for a certain welded assembly. They conducted a study on a total of four assembles. However, the data on samples 2 and 3 were corrupted. The records for sample 1 and sample 4 were 6 hours and 4.86 hours respectively. Determine the learning rate for this assembly using (a) doubling method and (b) formula method.


this is the whole question, nothing was added to it.
Answer #1

a)

Doubling Method -

Time Required for nth unit = T*L^x

Where,

T = Time Required for First unit

L = learning Rate

x = Number of time T will be doubled

Here, x will be equal to 2 as there will be twice doubling, 1 to 2 to 4

Time Required for 4th Sample = 6*(L)^2

4.86 = 6*L^2

L^2 = (4.86/6)

L = (4.86/6)^(1/2)

L = 0.9

Learning Rate = 90%

b)

Tn = T1*N^b

Where

T1 = Time required for first unit

N = Unit Number

b = ln(learning rate)/ln(2)

4.86 = 6*4^b

4^b = (4.86/6)

4^b = 0.81

b = ln(0.81)/ln(4)

b = -0.152

ln(Learning Rate)/ln(2) = -0.152

ln(Learning Rate) = -0.152*ln(2)

ln(Learning Rate) = -0.105358

Learning rate = e^-0.105358

Learning Rate = 0.90 or 90%

Hope I am able to solve your query. In case you find it useful, kindly provide positive feedback by doing Thumbs up.